OpenStudy (anonymous):

Solve each system of equations by substitution y = –2x – 5 y = 3x + 20

OpenStudy (anonymous):

substituting y, on 2nd eqn -2x-5=3x+20 5x=-25 x=-5 y=(-5*-2)-5=10-5=5